OpenStudy (anonymous):

hmm i need help badly find f [g(2)]

OpenStudy (anonymous):

For all 5 problems in this assignment, use the following functions: f(x) = x + 5 g(x) = x2 -4x + 5 h(x) = 3x2 - 2x + 5

OpenStudy (anonymous):

f[1]=6

OpenStudy (anonymous):

thats the answer?

OpenStudy (anonymous):

i got something totally different

OpenStudy (anonymous):

first find \(g(2)\) it is \(g(2)=2^2-4\times 2 +5=4-8+5=1\)

OpenStudy (anonymous):

then find \(f(1)\)

OpenStudy (anonymous):

how ?? @janessia

OpenStudy (anonymous):

what @moha_10 said

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

wait so wat she gave me is the answer?

OpenStudy (anonymous):

yes, it is \(f(g(2))=f(1)=6\)

OpenStudy (anonymous):

ohhhh ohk makes since when u put it like tht im kinda slow

OpenStudy (anonymous):

thank you guys

OpenStudy (anonymous):

it's okay ur welcome

OpenStudy (anonymous):

lol :)